OpenStudy (anonymous):

a field is to fertilized at a cost of $0.08 per square yard. The rectangular part of the field is 95 yards long and the diameter of each simicircle is 49 yards. find the cost of fertilizing the field use 3.14

OpenStudy (anonymous):

pi*r^2= are of a circle, two semicircles= 1 circle 3.14*(49/2)^2 ==1884.79 yd^2 area rectangle = length*width so 95*49=4655 add it all up.... 4655+1884.79=6539.79

OpenStudy (anonymous):

ummm lol thts not one of the answers.

OpenStudy (anonymous):

what does the field look like?

OpenStudy (anonymous):

i assumed it was a rectangle with two half circles at each end

OpenStudy (anonymous):

yea ur right lol